OpenStudy (anonymous):

the minute hand on the clock is 2m long. What is the mean angular speed?

OpenStudy (ujjwal):

the angular speed is equal to angular displacement divided by time taken.. for an angular displacement of 2 pi radians, it takes 60*60 sec.. so angular speed = 2pi/(60*60)= your answer.. the length of minute hand is not significant though..

OpenStudy (anonymous):

yeah , I did that, my ans is 1.7 x10^-3 rad/sec, which is wrong

OpenStudy (ujjwal):

How can it be wrong?? what is the correct answer then?

OpenStudy (anonymous):

what I have is 1x10^-3

OpenStudy (ujjwal):

they ate 0.7*10^-3... lol.

OpenStudy (anonymous):

haha, and now we're banging our heads cuz of this

OpenStudy (anonymous):

v=2pir/t r=2m t=60s so its 2*pi*2/60=pi/15

OpenStudy (anonymous):

v=omega*r omega=v/r=pi/15/2=0.1047

OpenStudy (anonymous):

@Mythias it will travel 2 pi r in 3600 seconds not 60sec, its the minute needle we're talking about.